# Disentangling Representations in RNNs through Multi-task Learning

### v1

@@ 5 Discussion
5.1 Implications for deep learning
Our work provides insight for understanding how and why “understanding” emerges from solving certain prediction tasks, which is particularly relevant to the study of foundation models [10, 31, 6]. Foundation models exhibit “understanding” of the world through semantically rich abstract representations that enable efficient downstream prediction. Our work suggests a profound connection between multi-task prediction and implicit knowledge of the underlying data generation process.
Our work provides a geometric starting point for understanding the correspondence between $\mathbf{x}^{*}$ and $\mathbf{Z}(t)$ in terms of mutual information $I(\mathbf{x}^{*};\mathbf{Z}(t))$ under various prediction tasks $\mathbf{y}(\mathbf{x}^{*})$. Relating the properties of prediction task $y$ with $\mathbf{x}^{*}$ and $\mathbf{Z}(t)$ has implications for optimal synthetic data generation and novel unsupervised learning objectives for foundation models. The formulation is fundamental, with clear informational and geometric constructions amenable to both theory and experiment.
We demonstrated that even simple binary multi-task classification forces an optimal estimator of $\mathbf{x}^{*}$ in $\mathbf{Z}(t)$. Binary classification is a low information objective, as it only requires the model to predict $N_{\textit{task}}$ bits that are deterministic functions of a continuous $D$-dimensional vector $\mathbf{x}^{*}$. The fact that an optimal estimator can be recovered under such conditions (i.e., $H(\mathbf{x}^{*})>>H(y_{1},\ldots,y_{N_{\textit{task}}})$) is remarkable. Higher resolution prediction targets may yield the same or better estimation of $\mathbf{x}^{*}$, yet we already proved that the multi-task binary case achieves optimal estimation. Therefore, predicting masked out parts of an image during self-supervised learning or next word prediction in LLMs, which are richer objectives, should perform at least as well as our experiments. In fact, evidence for the latter already exist at the form of abstract, monosemantic concepts in LLMs [70].
Finally, the present work emphasizes the importance of the diversity and quality of data, rather than quantity, for generalization (an argument also made by Arora and Goyal [3]). Indeed, our experiments show that few uniformly spaced classification boundaries outperform many that over-sample part of the state space, ignoring others (Figure 1f, Figure S9b,c). Thus, we hope that this work will serve as inspiration for developing efficient and generalizable deep learning systems.
@@ 5.2 Implications for neuroscience
An ongoing debate in the brain sciences is whether to solve tasks the brain learns abstracts representations, or simple input-to-output mappings. Here we show that training RNNs to multitask results in shared, disentangled representations of the latent variables, in the form of continuous attractors. In this multitask setting, one task acts as a regularizer for the others, by not letting the representation collapse, or overfit, to specific tasks [77]. In the brain, such pressure from multiple tasks could be imposed by thalamocortical loops for internal state-dependent action selection operating in parallel, while the integration occurs in corticostriatal circuits [63]. This framework is also consistent with proposed theories of parallel processing in cortical columns [27].
Our findings directly link to two important neuroscientific findings: spatial cognition and value-based decision-making. First, the tasks here bear close resemblance to path-integration, i.e. the ability of animals to navigate space only relying on their proprioceptive sense of linear and angular velocity [52, 12, 72]. In path-integration animals integrate velocity signals to get location, while here we integrate noisy evidence to get rid of the noise. We learn abstract representations in the form of a 2D "sheet" continuous attractor, while the computational substrate for path integration is a 2D toroidal attractor [25, 68] – not an abstract representation. The conditions under which a 2D sheet vs. toroidal continuous attractor is learned is a potential area of future research. Second, decision making experiments in monkeys result in a 2D abstract representation in the medial frontal cortex, which supports novel inferential decisions [7]. Likewise, context-dependent decision-making experiments in humans also resulted in orthogonal, abstract representations [24].
The workhorse model for computational neuroscience has been context-dependent computation, where tasks are carried out one at a time and task identity is cued to the RNN by a one-hot vector [48]. However, this approach can be algorithmically inefficient, scaling linearly with the number of tasks, and exponentially with input dimensionality $D$. That is because context-dependent computation utilizes different parts of the state space for different tasks, and the resulting representations collapse to what is minimally required for each task (see [48, 75] and Section A.4). This can be detrimental for brains, which need to pack a lot of computation within a large yet limited neural substrate. In contrast, abstract representations are general, compact [47], can be used for any downstream task involving the same variables, scale linearly with $D$, and readily emerge from relatively simple tasks.
5.3 Limitations
A limitation of the present work is that factorization is assumed. Yet not all problems are factorizable, or should be factorized. For instance, a more coarse-grained understanding of the world, that doesn’t disentangle all factors, might be more suitable in many cases, and that might be reflected in the nature of the tasks. Furthermore, we focus on canonical cognitive neuroscience tasks which are somewhat removed from standard ML benchmarks. Normally, disentanglement methods would be tested against a benchmark such as dSprites [50]; however to the best of our knowledge no such benchmark exists for dynamic tasks where evidence has to be aggregated over time. Future work could endeavor to apply our setting to ML tasks, like self-supervised image completion, and showcase the potential benefits of scaling up the number of masked out pixels in relation to the dimensionality of the latent space $D$. Furthermore, we assume that noise is Gaussian and IID, which might not be always the case; yet our theory accounts for other noise distributions as long as the posterior $P(\mathbf{x}^{*}|\mathbf{X}(1),\dots,\mathbf{X}(t))$ is non-zero everywhere and has a self-consistent/invariant maximum likelihood estimate over $\mathbf{x}^{*}$ (see Appendix B).
5.4 Conclusion
We focus on the problem of optimal estimation of a latent ground truth $\mathbf{x}^{*}$ from noisy, non-linear, high-dimensional versions of it. We provide theoretical guarantees that estimation succeeds when the amount of partial information about the truth available $N_{\textit{task}}$ scales linearly with the dimensionality of the latent space $D$. Our results hold for any injective mapping of latent variables to observables, which incorporates a vast range of transformations. They offer exciting implications for the manifold hypothesis [22], proving that the latent representations of optimal evidence aggregators observing high-dimensional world data inevitably estimate the structure of the low-dimensional manifold. Of particular interest is extending the theory to non-injective and non-deterministic observation maps $f$.
These findings shed light in the conditions under which biological and artificial systems alike develop representations that generalize well: they do so when there is enough pressure from many tasks that involve the same latent variables. Quoting Bengio et al. [4]: ’.. the most robust approach to feature learning is to disentangle as many factors as possible, discarding as little information about the data as is practical’. Apart from understanding learning in brains, we hope this work will inspire the development of deep learning systems with OOD generalization in mind.
@@ 5.4 Conclusion
PV would like to thank the Onassis Foundation and AR the NOMIS Foundation for funding. AB thanks the NIH PTQN program for funding. No competing interests to declare. We would like to thank Yisong Yue and Stefano Fusi for early discussions, and Aiden Rosebush for early discussions of proof methods.
@@ Appendix A Supplementary materials
A.1 Finding fixed points of the dynamics
To find the fixed points of the network dynamics after training, we follow a standard procedure outlined in Sussillo and Barak [69]. Specifically, we keep network weights fixed, provide no inputs to the network, and instead optimize over hidden activity. Specifically, we penalize any changes in the hidden activity, motivating the network to find stable states of the dynamics in the absence of input, i.e. attractors of the dynamics. This process finds all states of accumulated evidence that can be stored in this network as short-term memory. Network dynamics could then leverage these states to maintain and update the internal representation of the ground truth $\mathbf{x}^{*}$ on a single trial level, and drive downstream decisions.
@@ A.2 Quantification of generalization performance
A.3 Derivation of theoretical $r^{2}$ for optimal multi-task classifiers
@@
Remembering that $x^{*}_{i}\sim\textit{Uniform}(-0.5,0.5)$ it follows that $Var(x^{*}_{i})=\frac{2}{3}0.5^{3}$, and replacing $\sigma=0.2$ from Table S1 we arrive to $r^{2}=1-\frac{0.48}{T}$ for given trial duration $T$ which we compare to RNN OOD generalization performance in Figure S9a.
A.4 State-space efficiency of context-dependent computation
@@
Here we compare context-dependent decision making, where one task is performed at a time [48], to multitasking, in terms of state-space usage efficiency. Compared to the network in the main text (Figure 1c), the RNN now also receives a one-hot task rule vector indicating the current task, and it outputs the decision for that task only (Figure S1b). We have also omitted the non-linear encoder, making the tasks easier. We train the RNN for two tasks, one task at a time, in interleaved batches (Figure S1a). In one task, the RNN is required to decide which stream has more evidence, and at the other whether the sum of evidence across streams exceeds a certain decision threshold (here 0).
@@
Figure S1: Representations in an RNN trained in context-dependent decision making. (a) We trained RNNs for two classification tasks: two-alternative forced choice (where the decision boundary is the $x_{1}=x_{2}$ line) and evidence integration (corresponding to the $x_{1}=-x_{2}$ line). Each task corresponds to a different one-hot task rule vector. (b) Network architecture. In addition to the inputs in Figure 1c, the network also a one-hot vector indicating the current task. (c), Top 3 PCs of RNN activity example trials (40 in total). The task rule biases the network towards learning separate solutions in different parts of the state space for different tasks, in the form of separate line attractors; red x’s for two-alternative forced choice and orange x’s for evidence integration.
@@
We find that in this setting the network is not incentivized to learn abstract representations. Instead a separate line attractor is present in the learned dynamics for each task (red and orange x’s in Figure S1c; one of the line attractors is presumably tracking the difference of evidence (similar to Yang and Wang [75] but for independent evidence streams) and the other the sum of evidence. That is to say, the task rule biases the network to learn different computations in separate regions of state space, as in Mante et al. [48]. As a result, the two dimensional latent space has collapsed and cannot be decoded from network activity; therefore generalization to any task that involves these two variables is not possible.

### v2

@@ 6 Discussion
In this work, we proved that disentangled, generalizable representations must emerge in agents optimally solving multi-task evidence accumulation tasks canonical in the neuroscience literature. We also conducted experiments in a suite of autoregressive models (RNNs, LSTMs, transformers) which confirmed all of the main theoretical predictions. A key takeaway is that transformers more readily disentangle representations, which may explain their unique world understanding abilities.
6.1 Implications for representation learning
Topology-preserving representation learning
Our work has profound implications for learning representations that inherit the topological structure of the world. We prove this naturally happens as long as there are enough tasks to uniquely identify the location of $\mathbf{x}^{*}$. Crucially, the constraints from different tasks should be placed simultaneously on the representation, which explains why representations from context-dependent computation [49] are typically not disentangled.
Representational alignment across individuals
Our results provide a new perspective on the Platonic representation hypothesis [34], which suggests that the convergence in deep neural network representations is driven by a shared statistical model of reality, like Plato’s concept of an ideal “Platonic” reality. Theorem 3.1 suggests that the key factor driving convergence is the diversity and comprehensiveness of the tasks being learned. As long as individuals are faced with similar day-to-day tasks that collectively span the space of the underlying data representation, convergence to a shared, reality-aligned representation can occur. This could explain why for example modern LLMs come to encode high-level, human-interpretable concepts [71].
Manifold hypothesis
While our problem is framed in terms of arbitrary injective observation map $f$, the formulation encompasses many scenarios relevant to the manifold hypothesis [23]. The function $f$ can represent a smooth manifold embedded in a high dimensional space, directly modelling the manifold hypothesis of deep learning. In neuroscience, $f$ could be a non-linear encoding of stimuli in a neural population response, connecting our work to neural manifold research [44]. By developing and testing theoretical guarantees for the emergence of disentangled representations in this multi-task problem formulation, we provide insight on how neural networks can inherently discover and linearize low-dimensional manifolds within high-dimensional, non-linear observations, enhancing our understanding of how complex data structures are captured and represented in deep learning models and biological systems alike.
Interplay between number of tasks and fine-grainness of representations
Finally, the theorem and experimental results provided here are not a one-way-street from dimensionality $D$ of the latents to how many tasks $N_{\text{task}}$ are required to uncover them. Instead, there is a fundamental interplay between richness of tasks performed and detail of the representation learned. In a high-dimensional world, the richness of the tasks at hand directly affects the dimensionality $D$ of the latents that can be extracted, allowing for "ground truths" $\mathbf{x}^{*}$ at different levels of granularity to be explored. The richer the label information available, the more fine-grained the resulting world model will be.
6.2 Limitations
A limitation of the present work is that factorization is assumed. Yet not all problems are factorizable, or should be factorized. For instance, a more coarse-grained understanding of the world, that doesn’t disentangle all factors, might be more suitable in many cases, and that might be reflected in the nature of the tasks. Furthermore, we focus on canonical cognitive neuroscience tasks which are somewhat removed from standard ML benchmarks. Normally, disentanglement methods would be tested against a benchmark such as dSprites [52]; however to the best of our knowledge no such benchmark exists for sequential tasks where evidence has to be aggregated over time. Future work could endeavor to apply our setting to richer tasks, like extracting latent item attributes from item embeddings when sequential decisions are made in online retailer settings.
@@ 6.3 Implications for neuroscience
The brain encodes variables of interest in a disentangled format, in processes as disparate as memory [8], emotion [67], and decision making [6]. Furthermore, performance in tasks has been shown to degrade once abstract representations collapse [67], supporting their role in guiding generalizable behavior. Given our findings, and that the cortical architecture is uniquely suited for parallel processing [28], the cortex is a prime candidate area for the construction of disentangled world models. Another such area is the thalamus; it is posited that thalamocortical loops operate in parallel, and combined with internal state-dependent mechanisms lead to state-dependent action selection (e.g. prioritizing water when thirsty), while evidence integration occurs in corticostriatal circuits [65]. The representations discovered here (continuous attractors, CANs) have been widely found in the brain when solving similar tasks, highlighting their role as a general computational substrate for cognitive functions in the brain (for relations of our work to the neuroscience literature, see Section A.11). The algorithmic efficiency of multi-task learning compared to alternatives (“context-dependent computation”, Mante et al. [49], Section A.7), makes us think that it is no coincidence that the cortex can support parallel processing; all the pieces are there, and we feel that the brain has to leverage this feature to construct faithful models of the world, as it does.
@@
PV would like to thank the Onassis Foundation and AR the NOMIS Foundation for funding. AB thanks the NIH PTQN program for funding. No competing interests to declare. We would like to thank Yisong Yue for early discussions, Stefano Fusi for early feedback and Aiden Rosebush for discussions of proof methods.
@@ Appendix A Supplementary materials
A.1 Definitions
In the main text we used the terms "abstractness" and "disentanglement". Since there exists some ambivalence about their meaning in the literature, we would like to strictly define them here. We will be using definitions adapted from Ostojic and Fusi [59]:
  • Abstract representation of $x_{1},\dots,x_{n}$ represents each $x_{i}$ linearly and approximately mutually orthogonally. Thus, abstractness ensures a simple linear map can decode each $x_{i}$ regardless of variation in $x_{j\neq i}$.
  • Disentangled representations of $x_{1},\dots,x_{n}$ encode each $x_{i}$ orthogonally, without the necessity of linearity.
Note that under this definition, axis-alignment is not a requirement for disentanglement (also see Higgins et al. [31]). Our work suggests that the computer science and neuroscience communities should adopt this broader definition of disentanglement, because otherwise we might be missing cases where the factors are not axis-aligned, but they are still orthogonal and can still be isolated by a linear decoder. Our argument is that there is nothing special about individual factors being encoded by individual neurons. Rather, we think that allowing for mixed representations within the definition of disentanglement leads to a more holistic view of disentanglement. A contribution of this work, along with others [37], is to bring this argument to the forefront.
@@ A.2 Quantification of generalization performance
A.3 Estimation of angles between latent factors
To estimate the angles between latent factors in the representation, we obtain the normal vectors of the decoders $\mathbf{A}$ for each of the latents, and compute pairwise angles for all of them. To account for variability in the decoder fits, we repeat the decoder fit 5 times for each out-of-distribution region (see Section A.2 for details). We also repeat this process across 5 trained networks for each combination of $(N_{\text{task}},D)$, and report the 25, 50 (median) and 75 percentiles of all values for each $(N_{\text{task}},D)$ combination in Figure 3c and Figure S10.
A.4 Derivation of theoretical $r^{2}$ for optimal multi-task classifiers
@@
Remembering that $x^{*}_{i}\sim\textit{Uniform}(-0.5,0.5)$ it follows that $Var(x^{*}_{i})=\frac{2}{3}0.5^{3}$, and replacing $\sigma=0.2$ from Table S1 we arrive to $r^{2}=1-\frac{0.48}{T}$ for given trial duration $T$ which we compare to RNN OOD generalization performance in Figure 5a.
A.5 GPT-2 experiments
We train GPT-2 causal transformers with $d_{\textit{model}}=N_{\textit{neu}}=64$, $N_{\textit{layer}}=1$, $N_{\textit{head}}=8$ in the multi-task classification task of the main text. The networks receive continuous, noisy and non-linearly mapped inputs $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$, and should output multi-task classifications $\mathbf{y}(\mathbf{x}^{*})\in\{0,1\}^{N_{\textit{task}}}$. The output of the network is $\hat{\mathbf{Y}}(t):=g(\mathbf{Z}(t))$, where $\mathbf{Z}(t)$ is the last embedding of the sequence in the last layer and $g=\text{sigmoid}$. Since the input is continuous, we omit the tokenization and embedding steps, and project the input directly to the hidden state with a linear map. Furthermore, since the inputs are IID, we do not include positional encodings. The networks are trained with binary cross-entropy loss for $N_{\textit{batch}}=2*10^{4}$ batches, while the rest of the parameters are identical to the fixed RT networks of the main text (Table S1).
A.6 Finding fixed points and linearization of dynamics
To find approximate fixed points of RNN dynamics after training, we follow a standard procedure outlined in Sussillo and Barak [70]. Specifically, we keep network weights fixed, provide no inputs to the network, and instead optimize over hidden activity. Specifically, we penalize any changes in the hidden activity, motivating the network to find stable states of the dynamics in the absence of input, i.e. attractors of the dynamics. This process finds all states of accumulated evidence that can be stored in this network as short-term memory. Network dynamics could then leverage these states to maintain and update the internal representation of the ground truth $\mathbf{x}^{*}$ on a single trial level, and drive downstream decisions.
Then for every approximate fixed point $\mathbf{z}^{f}$, we linearize RNN dynamics around it and estimate the eigenmodes which describe how the system behaves in a small region $\delta\mathbf{z}$ around $\mathbf{z}^{f}$. Specifically, following Sussillo and Barak [70], Mante et al. [49] we take the difference system $\delta\mathbf{z}(t+t_{0})=\mathbf{z}(t+t_{0})-\mathbf{z}(t_{0})$ and linearize it, i.e.
$\dot{\delta\mathbf{z}}=\mathbf{F}^{\prime}(\mathbf{z}^{f})\,\delta\mathbf{z}$ (9)
where $\dot{\mathbf{z}}=\mathbf{F}(\mathbf{z})$ is the function describing the RNN dynamics and $\mathbf{M}\equiv\mathbf{F}^{\prime}(\mathbf{z}^{f})$ is its Jacobian at $\mathbf{z}^{f}$. To estimate $\mathbf{F}^{\prime}(\mathbf{z}^{f})$, we let network dynamics run in the absence of inputs for one time step $\Delta t$ starting from $\mathbf{z}^{f}$, i.e. $\delta\mathbf{z}(\Delta t)=\mathbf{z}(\Delta t)-\mathbf{z}^{f}$, and autodifferentiate $\delta\mathbf{z}(\Delta t)$. We then perform eigendecomposition of $\mathbf{M}$ and report the eigenvalues around each approximate fixed point. Eigenvalues near 00 indicate that the difference system $\delta\mathbf{z}(t)=\mathbf{z}(t)-\mathbf{z}^{f}$ changes slowly over time, i.e. they correspond to "slow" dimensions in network dynamics which can integrate inputs and maintain them over time (continuous attractors) [3, 49].
A.7 State-space efficiency of context-dependent computation
The workhorse model for computational neuroscience has been context-dependent computation, where tasks are carried out one at a time and task identity is cued to the RNN by a one-hot vector [49]. However, this approach can be algorithmically inefficient, because as we show here it scales linearly with the number of tasks $N_{\textit{task}}$, and exponentially with input dimensionality $D$. This is because context-dependent computation utilizes different parts of the state space for different tasks, and the resulting representations collapse to what is minimally required for each task (also see [49, 77]). This can be detrimental for brains, which need to pack a lot of computation within a large yet limited neural substrate. In contrast, abstract representations are general, compact [48], can be used for any downstream task involving the same variables, scale linearly with $D$, and as we show readily emerge from relatively simple tasks.
@@ A.7 State-space efficiency of context-dependent computation
To compare context-dependent decision making, where one task is performed at a time [49], to multitasking, in terms of state-space usage efficiency, we train RNNs to perform context-dependent decisions on the same tasks encountered in the main text. Compared to the network in the main text (Figure 2b), the RNN now also receives a one-hot task rule vector indicating the current task, and it outputs the decision for that task only (Figure S1b). We have also omitted the non-linear encoder, making the tasks easier. We train the RNN for two tasks, one task at a time, in interleaved batches (Figure S1a). In one task, the RNN is required to decide which stream has more evidence, and at the other whether the sum of evidence across streams exceeds a certain decision threshold (here 0).
@@
Figure S1: Representations in an RNN trained in context-dependent decision making. (a) We trained RNNs for two classification tasks: two-alternative forced choice (where the decision boundary is the $x_{1}=x_{2}$ line) and evidence integration (corresponding to the $x_{1}=-x_{2}$ line). Each task corresponds to a different one-hot task rule vector. (b) Network architecture. In addition to the inputs in Figure 2b, the network also a one-hot vector indicating the current task. (c), Top 3 PCs of RNN activity example trials (40 in total). The task rule biases the network towards learning separate solutions in different parts of the state space for different tasks, in the form of separate line attractors; red x’s for two-alternative forced choice and orange x’s for evidence integration.
@@
We find that in this setting the network is not incentivized to learn abstract representations. Instead a separate line attractor is present in the dynamics for each task (red and orange x’s in Figure S1c); one of them is presumably tracking the difference of evidence (similar to Yang and Wang [77] but for independent evidence streams) and the other the sum of evidence. That is to say, the task rule biases the network to learn different computations in separate regions of state space, as in Mante et al. [49]. As a result, the 2D latent space has collapsed and cannot be decoded from network activity; therefore generalization to any task that involves these two variables is not possible.
@@
A.8 Robustness to other noise distributions and correlated inputs
We here show that our setting is robust to Gaussian anisotropic and autocorrelated noise, and other asymmetric distributions of noise (Gumbel) whose CDF no longer matches the sigmoid functions in shape, with almost no drop in performance, and correlated inputs. This demonstrates that abstract representations are also learned outside of the specific assumptions made by our theory.
Starting from anisotropic noise, we observe that doubling the standard deviation of noise across one dimension ($\sigma=0.4$) does not result in a reduction in OOD generalization performance (median $r^{2}=0.96$ for $N_{\textit{task}}=24$, $D=2$). This is in line with our theory that can be extended to cover anisotropic noise (Lemma B.10). Non-IID noise should not be a problem either, since we are training our network for many samples and the effects of correlations will cancel out over long ensembles. Indeed, we find that including autocorrelated AR(1) noise with an AR coefficient of 0.7 results in only minor reduction in performance (median $r^{2}=0.95$ for $N_{\textit{task}}=24$, $D=2$).
We were also curious to see the impact of correlated inputs. A problem with high correlations is that they render parts of the state space virtually invisible to the network (Figure S2a). Surprisingly, OOD generalization performance is very weakly affected by input correlations, even though the state space is sampled uniformly in test (Figure S2b). The behavior is highly non-linear: performance is great until $\rho=0.97$, but for perfectly correlated inputs ($\rho=1$), the performance drop is sharp.
Figure S2: Disentanglement and factor correlations. (a) We introduce strong correlations in the latent factors, rendering parts of the state space virtually invisible to the network during pre-training (trained for a total of 24 classification tasks). (b) Despite that, generalization performance is excellent for correlations very close to 1. Once the factors are perfectly, performance drops significantly. This implies that the network can learn an abstract representation from correlated inputs, as long as there is some signal about the factors independently. This finding goes beyond [37] to show that the multi-task learning setting allows OOD generalization when the distribution during training the RNN itself is vastly different that the one during testing.
Finally, our theory pointed out sigmoid functions as a choice for activation function because of their close resemblance to the Gaussian CDF, resulting in the best OOD $r^{2}$. Still we find that for an asymmetric noise distribution (Gumbel) whose CDF does not match sigmoid functions well, there is only a slight drop in performance (median $r^{2}=0.95$ from $0.96$). Therefore the conditions for the activation function/CDF should be quite lax; any monotonic bijective function should work with small performance drop. This drop in performance is because the representation would be “stretched out” and “compressed” in a non-linear manner in regions where there is discrepancy between the noise CDF and the activation function. But this nonlinear squishing (determined by the term $\Phi^{-1}(g(\mathbf{Z}(t)))$) would be geometrically inoffensive — no cutting or gluing together would be required to map from $\mathbf{Z}(t)$ to a linear representation of $\mu(t)$. As a result, the representations would remain approximately linearly decodable.
A.9 Nonlinear classification boundaries and interleaved learning
In the main text we trained networks on linear classification boundaries. The tasks are still non-linear, since the encoder renders these boundaries non-linear to the network. However, there are cases where the latent factors themselves might need to be combined non-linearly, to make decisions. For instance, if the two factors represent the amount and probability of reward respectively, an agent needs to multiply the two and decide whether the expected value exceeds a certain (metabolic) cost $\gamma$ of performing an action to obtain said reward. Figure S3a shows the classification lines for the multiplicative task, where the network should decide whether the ground truth $\mathbf{x}^{*}$ lies above or below the curve $x_{1}\,x_{2}=\gamma$, for multiple values of $\gamma$. This family of tasks is not covered by Theorem B.6, because they violate the injectivity condition. Hence, we wondered how the representation would look like if the network was trained on both the linear and multiplicative boundaries, as animals do.
Figure S3: Interleaved learning of linear and non-linear boundaries. (a) Classification lines for the multiplicative task. There is a total of 48 classification lines, 12 per quadrant. (b) The network learns an abstract representation when trained for the linear and multiplicative boundaries in interleaved batches.
For that we perform interleaved training of both tasks (i.e. train in batches sampled from one of the tasks at a time), a setting where neural networks excel at, compared to humans who excel at blocked training, where tasks are learned sequentially (but see Flesch et al. [25]). Figure S3b shows that the network still learns a disentanlged, two-dimensional continuous attractor. OOD generalization for this network is excellent, and almost identical to ID performance (median $r^{2}=0.94,0.97$ respectively). Overall, we conclude that our framework extends to interleaved learning of a mixture of linear and nonlinear boundaries, which better reflects the challenges encountered by agents in the real world. Note that during interleaved training, linear and non-linear tasks are not performed simultaneously; yet they are in immediate succession which can also place pressure to the network to gradually learn representations that satisfy all tasks. The relation between multi-task and interleaved learning is a promising topic for future research.
A.10 Abstract representations are learned for a free reaction time, integrate to bound task
In the main text we trained networks to produce a response at the end of the trial. However, in many situations agents are free to make a decision whenever they are certain enough. Therefore, we here seek to extend our framework to free reaction time (RT) decisions. A canonical model accounting for choices and reaction times in humans and animals is the drift-diffusion model [40, 11]. It is composed of an accumulator that integrates noisy evidence over time, until a certain amount of certainty, represented by a bound, is reached, triggering a decision. In the linear classification task setting, the accumulated amount of evidence at time $t$ for a line with slope $\alpha$, $A_{\alpha}(t)$ is given by:
$A_{\alpha}(t)=A_{\alpha}(t-1)+X_{1}(t)-\alpha\,X_{2}(t)$ (10)
Figure S4: Free reaction time task. (a) Data generating process. Every classification line from Figure 2a now corresponds to an accumulator (see corresponding colors), and the desired output for the RNN is the accumulator values for the entire trial. The accumulator is quantized to integer values between ±$\pm 5$. (b) Representation for RNN trained on free reaction time task. The network learns a two-dimensional continuous attractor, similar to Figure 3d. A 3D rotating figure to better visualize this representation is provided in the Supplementary Material. (c) OOD generalization performance for the free reaction time (RT) task. Free RT outperforms fixed RT for a small number of tasks.
Intuitively, $A_{\alpha}(t)$ reflects the amount of confidence at time $t$ that the ground truth $\mathbf{x}^{*}$ lies above or below the classification line with slope $\alpha$. Essentially, the network has to explicitly report distance from the classification lines, not just in which side of the line $\mathbf{x}^{*}$ lies for that trial. We set the decision bound to ±$\pm 5$, and plot the accumulators $A_{\alpha}$ for all lines in Figure S4a. Note that once the bound is reached a decision is effectively made and $A_{\alpha}$ is kept constant. Also, instead of using continuous values, we quantize $A_{\alpha}$, because it is going to be used as target signal to train the network, and we do not want to introduce a strong inductive bias towards integrating the evidence streams.
We then train the RNN to reproduce confidence estimates from Figure S4a for the entire trial. Compared to previous experiments, the fixation input is no longer available to determine when to produce a decision. Instead, decisions evolve dynamically throughout the trial. We also use a MSE loss, change the activation function to $g=5\tanh$, and the Adam learning rate $\eta_{0}=3*10^{-3}$, but all other parameters remain the same as in the main text. To have a closer correspondence to the free RT experiments here, we also train the fixed RT task from the main text with MSE loss, symmetric labels $\mathbf{y}(\mathbf{x}^{*})\in\{-1,+1\}^{N_{\textit{task}}}$ and output non-linearity $g=\tanh$. We find that the change of objective and loss only has minor effects on generalization performance.
Figure S4b shows that in this setting the network still learns a two-dimensional continuous attractor of the latent space. Furthermore, the free RT outperforms the fixed RT network from the main text (Figure S4c) for a small number of tasks, since it is explicitly required to report distance from the classification lines. However, as our theory shows (Lemma B.3)) the fixed RT network is also implicitly reporting distance from the boundaries, when behaving like an optimal multi-task classifier, which explains the similar performance for a larger number of tasks. Overall, we showed that our setting accounts for naturalistic free RT decisions, and provides theoretical justification for the importance of confidence signals in the brain [66, 51].
A.11 Relation to neuroscience literature
An ongoing debate in the brain sciences is whether to solve tasks the brain learns abstracts representations, or simple input-output mappings. Here we show that training RNNs to multitask results in shared, disentangled representations of the latent variables, in the form of continuous attractors. In this multitask setting, one task acts as a regularizer for the others, by not letting the representation collapse, or overfit, to specific tasks [79].
Our findings directly link to two important neuroscientific findings: spatial cognition and value-based decision-making. First, the tasks here bear close resemblance to path-integration, i.e. the ability of animals to navigate space only relying on their proprioceptive sense of linear and angular velocity [54, 12, 73]. In path-integration animals integrate velocity signals to get location, while here we integrate noisy evidence to get rid of the noise. In path-integration, networks have to explicitly report distances, while in our setting distances are estimated implicitly (Lemma B.3)). We learn abstract representations in the form of a 2D "sheet" continuous attractor, while the computational substrate for path integration is a 2D toroidal attractor [26, 69] – not an abstract representation. The conditions under which a 2D sheet vs. toroidal continuous attractor is learned is a potential area of future research. Second, decision making experiments in monkeys result in a 2D abstract representation in the medial frontal cortex, which supports novel inferential decisions [6]. Likewise, context-dependent decision-making experiments in humans also resulted in orthogonal, abstract representations [25].